OpenStudy (anonymous):

For what value of "a" is f(x)= x^2-1, x=<3 2ax , x >3 continuous over all Real numbers?

OpenStudy (anonymous):

replace x by 3 in both expressions set them equal and solve for a

OpenStudy (anonymous):

\[f(3)=3^2-1=9-1=8=2a\times 3=6a\]

OpenStudy (anonymous):

\[6a=8\] \[a=\frac{8}{6}=\frac{4}{3}\]

OpenStudy (anonymous):

Brilliant, thank you very much!